David's Astronomy Pages
Notes - Session 1315 (2025-04-11)

 
Bullet Session Aims & Highlights
 - Observing Result
 - Night Summary Plot
 - Session Event Log
Bullet Operational Issues        
  - Critical Issues (0),  Major Issues (1),  Minor Issues (4),  Small Defects (15),  Continuous Improvement (20)
    (Note : Issues & Defects mostly relate to post-session program development) 
Bullet Images from 2025-04-11 >>         
2025-12-12 to 2015-05-22
Bullet Geomagnetic Station - Installation (2025-04-12 to 2025-04-14)
Bullet Geomagnetic Station - Commissioning (2025-04-14 to 2025-04-24)
   
Bullet Daytime Geomagnetic Storm (2025-04-16)
   
   

Session Aims & Highlights (2025-04-11)

Main aims

  1. Targets.  Acquire images of a selection of variable stars, nearby stars, comets & deep sky targets as allowed by sky conditions.
  2. C/2025 F2 (SWAN). Acquire further images of new comet C/2025 F2 (SWAN)

 Equipment & Software

Highlights

Notes

Summary Plots & Logs

Observing Plan
Image
  
Observing Result
Image
   
Image
 
Dome & Scope Slewing Performance
Image
  
Slew/Centering Performance
Image
  
Guiding Performance
Image
  
Sky Conditions (Locate Frames)
Image 
  
Night Sky Summary Plot
Top axis: Sky Brightness at Zenith (in ADU/s)
Lefthand axis: Local Time (hh LT). Righthand axis: Sun Altitude (degs)
Image   
  
Actual Weather vs Pre-Session Weather Forecast
Image
Image
  
Session Event Log
Time     Event Detail
21:21:19 Session Monitoring AutoStart monitoring for Live Session opportunity between 21:21 & 03:52
21:21:22   AutoStart Delayed Session delayed due to too few stars
21:28:37 Session AutoStarting Session autostarting (21:28)
21:29:00   CCDSoft Restarting CCDSoft being restarted (to set AutoSave No.)
21:29:28   Camera1 Connected SBIG Camera connected (Set point -20°C)
21:29:30 Session Created Session Created (Live, 2025-04-11 S01315, ImageSaveNum: 1315001)
21:29:35   Scope Switched On Telescope Power has been switched on via UPB Powerbox.
21:31:18   Services Started Observatory Services started
21:31:26 Observatory (Auto) Observatory placed in Fully-Automated Mode
21:31:28   AutoStart Delayed Session delayed due to too few stars
21:31:30 Session Pending Session pending (2025-04-11)
21:31:32 Session Initiating Session initiating (2025-04-11)
21:31:39   Camera1 Connected SBIG Camera connected (Set point -15°C)
21:31:41   Plan Requested Observing Plan requested from AstroPlan (1.43)
21:32:07   Plan Loaded Observing Plan loaded to queue (Plan ID: 1099)
21:32:48   Telescope Connected Telescope connected (TheSky6)
21:33:12 Session Equilibration Session ready to Open Dome
21:33:57   Dome Opened Dome opened (Opening time 45s, Zigbee 40s)
21:34:43 Session Running Session running
21:34:45   Queue Started Observing Queue started (23 targets selected)
21:34:48     Target Started (NrZen) Target started (Focus Field 10, HIP 49253)
21:35:02       Dome Unparked Dome unparked
21:39:37       Focusing Started-Foc1 Foc1 Focusing Started (TCF-S)
21:42:23       Focusing Completed Foc1 AutoFocus Completed (Profile No 1, wide)
21:44:45       Focusing Completed Foc1 AutoFocus Completed (Profile No 1)
21:44:47       Focusing Started-Foc2 Foc2 Focusing Started (Secondary Scope, using ShCap)
21:47:00       Focusing Completed Foc2 AutoFocus Completed (Profile No 2, wide)
21:48:49       Focusing Completed Foc2 AutoFocus Completed (Profile No 2)
21:49:08     Target Completed (NrZen) Target completed (Focus Field 10, HIP 49253)
21:49:13     Target Started (1/23) Target started (1/23, NGC 2890 w/SN2023xnl)
21:56:14       Focusing Started-Foc1 Foc1 Focusing Started (TCF-S)
21:58:32       Focusing Completed Foc1 AutoFocus Completed (Profile No 3)
22:12:06     Target Completed Target partially completed (1/23, NGC 2890 w/SN2023xnl)
22:12:10     Target Started (2/23) Target started (2/23, AT2024aasb)
22:16:56       Focusing Started-Foc1 Foc1 Focusing Started (TCF-S)
22:19:34       Focusing Completed Foc1 AutoFocus Completed (Profile No 4)
22:35:46     Target Completed Target completed (2/23, AT2024aasb)
22:35:50     Target Started (3/23) Target started (3/23, Vega)
22:42:07     Target Completed Target completed (3/23, Vega)
22:42:11     Target Started (4/23) Target started (4/23, SN 2023nlk)
22:46:07       Focusing Skipped Foc1 focusing skipped - star is too dim (TCF-S)
23:02:31     Target Completed Target completed (4/23, SN 2023nlk)
23:02:35     Target Started (5/23) Target started (5/23, 29P/Schwassmann-Wachmann)
23:07:09       Focusing Started-Foc1 Foc1 Focusing Started (TCF-S)
23:09:38       Focusing Completed Foc1 AutoFocus Completed (Profile No 5)
23:25:31     Target Completed Target completed (5/23, 29P/Schwassmann-Wachmann)
23:25:35     Target Started (6/23) Target started (6/23, T CrB)
23:30:03       Focusing Started-Foc1 Foc1 Focusing Started (TCF-S)
23:32:11       Focusing Completed Foc1 AutoFocus Completed (Profile No 6)
23:37:21     Target Completed Target completed (6/23, T CrB)
23:37:25     Target Started (7/23) Target started (7/23, UGC 10058 w/SN2025gvs)
23:40:56       Focusing Started-Foc1 Foc1 Focusing Started (TCF-S)
23:43:32       Focusing Completed Foc1 AutoFocus Completed (Profile No 7)
00:00:00     Target Completed Target completed (7/23, UGC 10058 w/SN2025gvs)
00:00:04     Target Started (8/23) Target started (8/23, NGC 3277 w/SN2025coe)
00:20:05     Target Completed Target completed (8/23, NGC 3277 w/SN2025coe)
00:20:10     Target Started (9/23) Target started (9/23, AT2025gcz)
00:25:56       Focusing Skipped Foc1 focusing skipped - star is lost (TCF-S)
00:43:05     Target Completed Target completed (9/23, AT2025gcz)
00:43:09     Target Started (10/24) Target started (10/24, NGC 5957 w/SN2025fvw, ToO)
00:47:41       Focusing Started-Foc1 Foc1 Focusing Started (TCF-S)
00:50:34       Focusing Completed Foc1 AutoFocus Completed (Profile No 8)
01:06:58     Target Completed Target completed (10/24, NGC 5957 w/SN2025fvw, ToO)
01:07:00     Target Missed (11/24) Target's time slot was missed (11/24, NGC 2407 w/SN2025gmc)
01:07:02     Target Started (12/24) Target started (12/24, KIC 8462852)
01:14:14     Target Completed Target completed (12/24, KIC 8462852)
01:14:18     Target Started (13/24) Target started (13/24, AT2022abs (CV))
01:17:54       Focusing Started-Foc1 Foc1 Focusing Started (TCF-S)
01:20:23       Focusing Completed Foc1 AutoFocus Completed (Profile No 9)
01:26:53     Target Completed Target completed (13/24, AT2022abs (CV))
01:26:57     Target Started (14/24) Target started (14/24, LEDA 1776479 w/SN2025gcv)
01:30:45       Focusing Skipped Foc1 focusing skipped - star is lost (TCF-S)
01:46:52     Target Completed Target completed (14/24, LEDA 1776479 w/SN2025gcv)
01:47:24     Target Started (15/24) Target started (15/24, AT2024xpy)
01:51:17       Focusing Started-Foc1 Foc1 Focusing Started (TCF-S)
01:54:25       Focusing Completed Foc1 AutoFocus Completed (Profile No 10)
02:10:55     Target Completed Target completed (15/24, AT2024xpy)
02:10:59     Target Started (16/24) Target started (16/24, AT2022cya (CV))
02:15:18       Focusing Started-Foc1 Foc1 Focusing Started (TCF-S)
02:18:11       Focusing Completed Foc1 AutoFocus Completed (Profile No 11)
02:24:16     Target Completed Target completed (16/24, AT2022cya (CV))
02:24:20     Target Started (17/24) Target started (17/24, AT2025fwb)
02:27:37       Focusing Started-Foc1 Foc1 Focusing Started (TCF-S)
02:30:06       Focusing Completed Foc1 AutoFocus Completed (Profile No 12)
02:46:46     Target Completed Target completed (17/24, AT2025fwb)
02:46:51     Target Started (18/24) Target started (18/24, AT2023hru (CV))
02:49:55       Focusing Started-Foc1 Foc1 Focusing Started (TCF-S)
02:52:16       Focusing Completed Foc1 AutoFocus Completed (Profile No 13)
03:08:33     Target Completed Target completed (18/24, AT2023hru (CV))
03:08:37     Target Started (19/24) Target started (19/24, AT2023ij (CV))
03:12:30       Focusing Started-Foc1 Foc1 Focusing Started (TCF-S)
03:15:09       Focusing Completed Foc1 AutoFocus Completed (Profile No 14)
03:21:25     Target Completed Target completed (19/24, AT2023ij (CV))
03:21:29     Target Started (20/24) Target started (20/24, AT2024slv (CV))
03:24:52       Focusing Started-Foc1 Foc1 Focusing Started (TCF-S)
03:27:30       Focusing Completed Foc1 AutoFocus Completed (Profile No 15)
03:33:35     Target Completed Target completed (20/24, AT2024slv (CV))
03:33:39     Target Started (21/24) Target started (21/24, GCVS BL Lac)
03:36:44       Focusing Started-Foc1 Foc1 Focusing Started (TCF-S)
03:39:11       Focusing Completed Foc1 AutoFocus Completed (Profile No 16)
03:43:47     Target Completed Target completed (21/24, GCVS BL Lac)
03:43:51     Target Started (22/24) Target started (22/24, AT2023ygv)
03:47:40       Focusing Started-Foc1 Foc1 Focusing Started (TCF-S)
03:50:00       Focusing Completed Foc1 AutoFocus Completed (Profile No 17)
04:06:52     Target Completed Target completed (22/24, AT2023ygv)
04:06:56     Target Started (23/24) Target started (23/24, AT2022pxo (CV))
04:15:51     Target Completed Target completed (23/24, AT2022pxo (CV))
04:15:55     Target Started (24/25) Target started (24/25, C/2025 F2 (SWAN), ToO)
04:20:00       Focusing Started-Foc1 Foc1 Focusing Started (TCF-S)
04:22:33       Focusing Completed Foc1 AutoFocus Completed (Profile No 18)
04:33:14     Target Completed Target completed (24/25, C/2025 F2 (SWAN), ToO)
04:33:18     Target Started (25/25) Target started (25/25, UGC 12293 w/AT2025gyi)
04:49:20     Target Completed Target partially completed (25/25, UGC 12293 w/AT2025gyi)
04:49:22   Queue Completed Job Queue completed
04:49:25 Session Closing Session closing
04:50:20   Dome Closed Dome closed (Closing time 50s, Zigbee 44s)
04:50:38   Dome Parked Dome parked (parking time 10s), Az: 90.0 deg
04:51:27   Telescope Parked Telescope parked (parking time 45s total)
04:51:51   Telescope State Handbox reads 'Scope parked Turn scope off.'
04:52:47   Telescope Switched Off Telescope Power has been switched off via UPB Switch.
04:54:12   Services Stopped Night Services stopped
04:54:15 Session Finishing Session Finishing started (Create Fits Summary, Transfer Files)
04:54:42     Dome (Find Park) Find Park started (Search Az 88.0 to 92.0°, Step 0.2°, Narrow)
04:55:27     Dome Find Park found (Best Park Az 90.4°)
04:55:34     Dome Dome has been synced at park position (Az 90.0°, Adjustment: -0.4°)
04:55:36   Dome Parked Dome parked
04:57:21   Dome Charging Dome is parked and re-charging Ok (18mA)
04:57:23 Session Finished Session Finished
 
Session Alerts & Alarms
Time     Type       Name Detail
04:57:21 Green Clear Dome Charging Dome is parked and re-charging Ok (18mA)
 

Back to Top


Operational Issues (2025-04-11, S1315)

[ Prev | Next ]

Critical Issues

Major Issues

Minor Issues

Small Defects

Continuous Improvement

[ Prev | Next ]


Fig 1.  Magnetogram showing anomalous shifts in By values (from Y Axis Sensor)

Image

Fig 2. Comparison of magnetometer noise levels 2025-04-08 (top)  & 2025-04-14 (bottom)

Fig 2a. The top graph (indoor testing, un-assembled) shows relatively little noise over the bulk of the 12 hour run
 (the slightly higher noise levels before 00:45 are very likely due to computer use & small movements on the workbench).

Image

Fig 2b. The bottom graph (outdoor testing, assembled unit, operating downhole but not yet buried) shows significantly higher noise levels over the 4 hour run.  

Both graphs are shown with the same Y axis range (500nT).

Image

Fig 3.  Magnetometer Plot from 2025-04-16 showing observed shifts in sensor readings

Plot shows an anomalous -283 nT shift in Y reading at 11:50 UT proceeded by an anomalouus +1400nT spike.

Plot shows shifts in X, Y & Z values coincident with turning off power to the Temperature Sensor.
Shifts amounts were  -716nT,  -548nT & -454 nT respectively.

The plot also shows geomagenetic activity in X sensor which corresponds with a storm as recorded by Scandinavia Magnetometers

Image


Fig 4.  A later Magnetometer Plot from 2025-04-16 / 2025-04-17 showing continued anomalous shifts in sensor readings

Although the Temperature probe remained turned off (which is at least in part implicated in Shift isssses there was a further anomalous shift in Y Axis Readings at 19:36 UT in the order of +1090 nT, and a corresponding through smaller shift in Z Axis reading of +120 nT, and a questionable shift in X axis readings of +50nT.

Image


Fig 5.  SAM View Error Dialog associated with program crashes after restarting program (2025-04-17).


Image

 

Fig 6. Series of Red/Green Alerts on 2025-04-17 22:30 to 22:31

Image

[ Prev | Next ]

Back to Top


2025-04-12 to 2025-04-19


Geomagnetic Station - Installation (2025-04-12 to 2025-04-14)

A Geomagnetic Monitoring Station is being added to the Observatory in Spring 2025 for monitoring changes in the earth's magnetic field and to supplement and build upon the Observatory's existing Aurora monitoring that use a Black/White AllSky Camera and new NorthCam Colour Camera.

The following notes describe the Installation Phase (Preparation & Installation)

Other Notes :
 | Review | Site Prep | Assembly Design | Parts Assembly | Testing | Final Assembly | Installation | Commissioning |

Installation Preparation (2025-04-12)

- External surface of pipe joints and the cable inlet/filler have now been double coated with PVC solvent glue.

- Cables have been bound into a cable bundle, and threaded through the flexible pipe.  There was too much friction/ bend to allow the cable to be passed through the entire 10.5m flexible cable in one go and it decided to cut the flexible pipe into 4 pieces to allow the cable bundle to be successfully passed through.

- The middle shelf in the Observatory Computing Bay has been cleared in preparation for placing the SAM III unit on. Lower shelf is confirmed to be loosely position which will allow it to be temporarily removed for drilling hole in floor of the Observatory Computing Bay for pulling through the cable bundle from the Magnetometer Sensor fixture.

Sensor Fixture completed and moved outside
Cables emerging from the pipework fixture are
Blue:  X Axis Cable (cat6), Red: Y Axis Cable (cat6)
Green:  Z Axis Cable (cat6), Yellow:  Temperature
    Sensor Fixture - Alternate View
Image   Image
      
Measuring & laying out flexible corrugated pipe
in preparation for taking cables back to Observatory
    Cables formed into a bundle
Image   Image
     
Cables formed into a bundle (foreground)
 in preparation for passing
through flexible corrugated pipe (background)
    Cable Bundle passed through Corrugated Pipe
There was too much friction / cable bend to pass through
 the 10.5m corrugated pipe in one go, and it was cut into 4 pieces
Image   Image
     

Installation Day 1 (2025-04-13)

- Hole has been marginally deepened (in the area of the Z Axis Sensor) to allow the entire Sensor Fixture to be placed as deeply as possible.
 
- A trench between the Sensor Hole and the Observatory has been dug for taking the flexible corregated pipe which contains the cable bundle from Magnetometer Sensor fixture.

- Flexible Pipe / Cable Bundle have been laid back to the Observatory, and cable bundle passed through hole into the Observatory.

- Flexible Pipe backed away from Observatory, and breaks between Pipe Sections joined with tape.

- Trench back filled and lawn repaired.

Trench dug back to Observatory
7cm wide & 20-25cm deep, 10m long
    Trench dug to below Observatory Computing By
Image   Image
      
Cable Bundle / Corrugated Pipe
ready be installed into trench
    Cable Bundle / Corrugated Pipe in trench
Image   Image
     
Cable Bundle / Corrugated Pipe
entering Sensor Hole
    Lawn after filling trench
and replacing turf
Image   Image
     

- Ardunio connected to Lindy USB Hub (in Observatory Computing Bay) - Appears as COM8 in DeviceManager

- AstroMag modified to use COM8 as default port for Temperature Data and issued as AstroMag 1.2.1

- Sabrent USB-Serial Convertor Driver installed on Observatory Computer

C:/Data/Computer/Home Computers/SABRENT USB-Serial Converter Driver/Windows/PL23XX_Prolific_DriverInstaller_v208/PL23XX-M_LogoDriver_Setup_v208_20211221.exe

- SAM III connected - Appears as COM9 in DeviceManager

Installation Day 2 (2025-04-14)

- Sensor Cables tied into SAM III Unit and Ardunio Project Box, using existing wire ends
(cables still need to be trimmed back to an optimal and equal length)

- Confirmed that signals are being received, and that SAM_VIEW and AstroMag are working

- Sensor Fixture has been installed in the Hole, oriented N-S , E-W & Vertical, half buried
- Last 1m of cable bundle has been buried at depth in the hole
- Hole back filled and lawn repaired
- Sensor Hole geolocated to allow the Sensor Fixture to be recovered in future, should it be required

- Sensor signals monitored for stabilisation

Sensor Fixture sitting at bottom of hole
Picture taken prior to alignment
    Sensor Fixture at bottom of hole (detail)
shows 1m length of cable bundle that will be buried
 at same depth as the sensor
also shows N-S axis (blue), E-W Axis (red) and
Vertical Axis (green, at bottom of hole)
Image   Image
      
Setting up North-South Survey Line
This was done by creating a line parallel to
Observatory's North-South Alignment
Picture is looking Due North along string line
(see Observatory N-S Alignment, 2018)
    Sensor Unit Positioned according to Survey Line
Unit positioned vertically, centred on the
North-South Survey String, and with X Axis aligned
with the North-South Survey String
4 canes pushed into the bottom of hole hold the
Sensor Unit in position pending burial
Image   Image
     
Sensor Unit Alignment
Due to location of sensor at bottom of hole
its alignment had to be done by eye.
    Sensor Unit in Finalised Position
ready for burial
Image   Image
     
Sensor Unit being buried
Z & Y Sensor Axes have been buried at this point
    Sensor Unit being buried
All 3 sensor axes and cable inlet have been
buried at this point. Just the top of the sensor unit is visible
Hole was subsequently back filled up to the level of the lawn
Image   Image
     

Update 2025-04-15 (Analysis / Potential Calibration)

Magnetograms from 1st day (2025-04-14) plotted and analysed.

First plot shows readings from the morning and early afternoon from time that Sam III unit was powered up with the Sensor Fixture positioned in the 1.2m deep hole, but not yet fully aligned and not yet buried.  Data trends are consistent with a negative relationship to temperature. Data is quite noisey.  Would there be a benefit in using moving averages ? (or average the 10s samples to create a 1 min binned average).

Image


After this the Sensor Unit was fully aligned and the hole backfilled with soil, during which readings were highly unstable and are not shown.

Other Station Notes :
 | Review | Site Prep | Assembly Design | Parts Assembly | Testing | Final Assembly | Installation | Commissioning |


Back to Top


Geomagnetic Station - Commissioning (2025-04-14 to 2025-04-22)

A Geomagnetic Monitoring Station is being added to the Observatory in Spring 2025 for monitoring changes in the earth's magnetic field and to supplement and build upon the Observatory's existing Aurora monitoring using AllSky Camera and new NorthCam Camera.

The following notes describe the Commissioning Phase.

Other Notes :
 | Review | SAM III | Site Prep | Assembly Design | Parts Assembly |  Testing | Final Assembly | Installation | Commissioning | Remedial Work |

Update 2025-04-15 (Analysis / Potential Calibration)

Magnetograms from 1st day (2025-04-14) plotted and analysed.

First plot shows readings from the morning and early afternoon from time that Sam III unit was powered up with the Sensor Fixture positioned in the 1.2m deep hole, but not yet fully aligned and not yet buried.  Data trends are consistent with a negative relationship to temperature.

Image

Data is quite noisey.  Would there be a benefit in using moving averages ? (or average the 10s samples to create a 1 min binned average).

Image
Average 10 min noise = 21.2 nT (range 14.3 to 40.9 nT)

The chart below shows the data with a 5 minute moving average

Image


After this the Sensor Unit was fully aligned and the hole backfilled with soil, during which readings were highly unstable and are not shown.

Once buried the reading were renormalised to 0 nT, and charting continued.

Image

      Y Z
Expected Values     16467 -100 47326
           
Prior Fixes       -833  +13584  +11815 
Currently used offsets      -82833  -68416  -70185 
           
Current Av Values      21540  13280  45030 
Additional Fixes     -5083  -13380  +2296 
           
Proposed New offsets      -87916  -81796  -67889 

Update 2025-04-16 (Analysis & Diagnostics) 

- Some issues occuring with AstroMag encountering with exceptions which causes SamMonitor loop to exit, which stops data recording.
Fixed in AstroMag 1.2.5.

- Spiky temperature values that started the day before was in partly due to temperature values being sampled every 60s for graphing. A more valid and smoother temperature is producing by averaging  temperature values (that are read every1s) over a 60s window.

- Anomalous jumps in Y reading are still occuring (eg. an anomalous -283 nT shift in Y reading at  2025-04-16 11:50 UT that was proceeded by an anomalouus +1400nT spike).  This is ranked a Major Issue.

- Began some diagnostics tests that aim to understand high noise levels, but might cast some light on the Sensor Shift issue.  The first test which was to turn-off the power to the Temperature Sensor at 13:23 UT which showed an immediate and significant downward shift in all 3 sensor reading , with X,Y, Z readings shifting by  -716nT,  -548nT & -454 nT respectively.

Image

- It was hoped that running the station with the Temperature Sensor turned off would fix the issue with anomalous shifts in Y reading with associated smaller shifts in X and/or Z reading, however after around 5 hours the problems with shifts in Y reading resumed again.

Image

Image



Update 2025-04-17 (Further Diagnostic Testing) 

- As part of continued testing SAM III Unit was turned off in order to disconnect the Y Magnetometer and to see if this would eliminate anomalous shifts in readings. Obviously the Y readings (which have shown the biggest shifts) won't be now available, but it is hoped to tell if smaller shifts in X & Z readings continue or not.  After a delay of 15 minutes of so the SAM III unit was restarted and data log resumed. Piecing together data it would seem that after turning on the X and Z Magnetometers (without the Y Magnetometer running) saw an upward shift in X and Z readings of +400 and +100 respectively.   The net shift gradually reduced to +350 and +50 after 1 hour.

Image

Image

Update 2025-04-18 (Further Observation) 

- Some 25 hours of observations have been completed with the Y Axis Magnetometer turned-off and there have been no anomalous shifts during this time

- Data Readings are still rather noisy and with, as yet, no definate explanation or cure.

Image

- The noise can be cleaned up using a 7 min moving average (averaging data +/- 3 mins of the plotted time point), and this highlights the key trends in the data which are of geomagnetic origin.

Image

- The geomagnetic origin is highlighted by the following comparison of the X Axis magnetometer readings and a stackplot of Scandinavian magnetograms which covers 2025-04-17 12:00 UT to 2025-05-18 12:00 UT.  The X Axis Magnetometer appears to show deviations that are twice as large as those shown by dob-sol-kar stations.  This isn't understood. Is the scale bar on the mkstackplot being misunderstood ?

Image

- Downhole Temperature Monitoring has been turned back on at 13:20 (12:20 UT) in order to understand its effect on X & Z magnetometers, beyond the expected initial shift in readings. Provide the shift is remains static it is manageable.

The magnetogram (below) shows there was a +370 nT shift in X Axis Readings and a +498 nT shift in Z Axis Readings when the Temperature Sensor was turned on.  These compare with the earlier shifts of  -716nT and -454 nT respectively when the Temperature Sensor was turned off two days earlier (2025-04-16 14:23 (13:23 UT).)

Image

During the first 6 hours with Temperature Sensor running there have been no anomalous shifts in X & Z readings
This would seem to point to the Y Axis Magnetometer as being the cause of the random large shifts in magnetometer readings.

Comparisons between X Axis Magnetomteter readings with a Stackplot of Scandinavian Magnetometer Data

1)  24h chart beginning 2025-04-18 01:00 UTImage

2)  24h chart beginning 2025-04-19 00:00 UT Image

3) 24h chart beginning at 2025-04-19 05:00 UT
Image 

4) 24h chart beginning at 2025-04-19 14:00 UT
Image

 

5) 24h chart beginning at 2025-04-19 14:00 UT
Image

 

6) 24h chart beginning at 2025-04-21 00:00 UT.
Image

Latest Magnetometer Data is now being displayed on Aurora Watch Page (Magnetometer Station section)

 | Review | SAM III | Site Prep | Assembly Design | Parts Assembly |  Testing | Final Assembly | Installation | Commissioning | Remedial Work |
 

A request to access/download ASCII data from Scandinavian/Norwegian Magnetometers has been sent to Magnar Gullikstad Johnsen at Tromsø Geophysical Observatory (TGO).

Magnetogram Plots for 2025-04-21

Comparison with magnetograms from Norway

Image

Image

Image

Activity Plot  (development in progress)


Image

When completed this will produce a chart based on that used by SAMNET's Activity Plot (like this one showing Sumburgh Activity Plot)

Noise Tests (2025-04-25)

Began some tests to see what might be contributing to the apparently elevated noise levels in the SAM magnetometer readings.
(remember the need to change only one thing at a time, and allow ample time to see/understand the effect of change, if any).

a) Temperature Sensor Test (previous test) - switched off USB power to Temperature Sensor. 
- no noticeable decrease in noise, but did cause and immediate shift in X, Y Z readings (X shift -716nT,  Y-548nT,  Z shift -454 nT)
- again no noticeable change in noise level when power to Temperature was later restored.

b) Power Supply Test 12:13 (11:13 UT) - switched the power supply to SAM III Unit from Observatory Bay / UPS Powered to the Observatory's main power panel

- Possible marked 50nT reduction in Z Axis Reading associated with change in power supply, but the reduction could be geomagnetic ?
  (latest magnetograms from Scandinavia (sol, bfe etc indicate that there has been no marked reduction in Z component)

- Possible 25nT reduction in X Axis Reading associated with change in power supply, it could be seem as either be geomagnetic and part of a gradual trend in reducing X reading since 08:00 UT, or a change that is not consistent with latest magnetograms from bfe and sol where H readings are current going up  (11:00 to 12:00 UT).

- It seems that a power supply change affects the X & Y base readings. This is supported by the shift in Z Axis readings when Power Reset test was conducted at 16:18 (15:18).

- Power Supply was switched back to Observatory Computing Bay / UPS Supported Supply at 20:17 (19:17 UT).
  No major change or shift in X and Z readings.  However the Z readings could be interpreted to have shifted by 25 nT. This is much less that the earlier shifts that were associated with supply change at 12:13 and power reset at 16:18

c) Power Reset Test 16:18 (15:18 UT) Power to SAM III Unit turned off for 1-2 minutes to understand how SAM View Responds

- SAM View not affected by powering off the SAM III Unit, though obviously it stops collecting data.   (Last Record 15:18:43:)
- SAM View not unduly affected by re-powering the SAM III Units, data collection recommences again (First New record  15:20:13)

X Axis reading seem to have been unaffected by the powet reset but the Z Axis reding appear to have reduced by 70 nT from  15:14 to 15:27. Whilst it was initially unclear whether this was caused by the SAM Power Reset or by turning off the Pulsar Dome Controller, further tests involving turning Power Controller off/on shows that Power Reset was responsible for the shift in X Axis Readings.

Unless there is some other significant reason for not using power from the Observatory Computing Bay it would seem best to power the SAM III unit from the UPS Supported Power Supply in the Obsrevatory Bay, so that the SAM III Unit and Magnetometers won't be reset due to any short power outage, which can produce both of data outage for the duration of the power outage and affect magnetometer readings when power is restored.

The following chart illustrates sensitivity of Z Axis Magnetometer to Power Resets where is power to SAM III unit is either changed to a different A/C plug or is simply reset.

Image

d) Pulsar Dome Test 16:23 (15:23) - Pulsar Dome Controller was turned off at 16:23 so that it's Induction Charger isn't trickle charging the Shutter Battery (a possible source of high frequency noise ?)

The test was carried out too close to the Power Reset Test to be diagnositic in itself.

Z Axis values reduced by 70 nT from  15:14 to 15:27, but it is unclear whether this was caused by the SAM Power Reset or by turning off the Pulsar Dome Controller.

- Dome Controller turned on at 18:00. There was no change in X or Z Readings. 
- Dome Controller turned off at 18:29. There was no change in X or Z Readings.
- Dome Controller turned on at 19:25. There was no change in X or Z Readings.

Pulsar Dome doesn't seem to have any affect on SAM readings or noise levels.

e) Single Axis Test 2025-04-26, 09:00 (08:00 UT) :  Sam III unit was turned off, and the Z Axis wires disconnected from the 10 pin terminal , leaving just the X Axis wires connected.   Turned SAM III unit back on, and monitored with SAM View & AstroMag.

- X Axis readings shifted by +675 nT, noise reduced to very low levels

See picture below :  which visually compares the noise in X Axis magnetometer readings with/without active Z Axis magnetometer (note : readings shifted to a common base for easier comparison).

Image

This shows that the observed elevayed noise issue is not related to the outdoor location (ie not related to Ground, Observatory or 3rd party related  interference), but is due to cross-talk between the different cables running closely in parallel for 10-12m.    This suggests that Shielded Cat6 cable should have been used instead of the unshielded CAT6 cable used.

f) Other Tests :  - The plan to make further tests (including a test to turn off householed Internet Over Powerline) have been suspended for time being in view of the result of the Single Axis test.   (Although the Observatory no longer uses Internet over Powerline ,  IOP signals might still be present in the power supply going to the Observatory and affecting the power supply to the SAM III Unit or producing ground current fluctuations.)

Y Axis Test (2025-04-26)

A test was begun today13:26 (12:26 UT) with all 3 magnetometers connected to the SAM III unit. This includes the Y Axis magnetometer which has been purposedly disconnected on 2025-04-17 09:00 due to bad spikes & shifts  in Y readings on 2025-04-16 & early 2025-04-17.   The new test was designed to see if the earlier issue with Y Axis Magnetometer is still present or has magically gone away.  This is key test in order to decide whether to trigger the 'Y Axis remedial plan' .

During the test Downhole Temperature increased from 8.70°C at 13:35 (12:35 UT) to 8.90°C at 14:35 (13.35).
which would be expected to lead to a decrease in magnetometer reading of 20 to 30nT.   (possiblly 50-100 nT for the Z Axis (original Y Axis)  ?)

- Reading still stabilizing at 14:35 (13:35 UT).
- Readings stabilised at 15:00 (14:00 UT)
- Offsets changed in AstroMag to X: -4939, Y: -12740, Z: 2605

- No problem with Y Axis reading as of 01:00 (00:00 UT) - 11 hours runtime.

Image

2 axes - Average 10 min noise level  : 2025-04-26 01:00 to 04:00   = 13.2 nT (1 SD),    Y --,   Z 17.3 nT
1 axis -  Average 10 min noise level  : 2025-04-26 09:00 to 12:00   = 2.1 nT (1 SD),     Y --,  Z --
3 axes - Average 10 min noise level  : 2025-04-26 15:00 to 18:00 (stable, 3 axes)  = 15.6 nT (1 SD),  Y 10.4 nT,  Z  20.9 nT
3 axes - Average 10 min noise level  : 2025-04-26 18:00 to 21:00 (stable, 3 axes)  = 16.2 nT (1 SD),  Y 15.5 nT,  Z  29.2 nT

However a significant spike occured over a 20s period at 13:25 (12:25) (after a prior 24h flawless period) which resulted in Y values being shifted by +40 to +50nT    X axis reading saw same spike and may have ended with a slight negative shift but this isn't definitive as readings were already on a downward trajectory. Z Axis readings were unaffected.

Image

Average 10 min noise level  : 2025-04-26 19:00 to 22:00 (stable, all 3 axes)  = 16.0 nT (1 SD),  Y 16.9 nT,  Z  31.3 nT
Average 10 min noise level  : 2025-04-27 00:00 to 03:00 (stable, all 3 axes)  = 16.7 nT (1 SD),  Y 12.0 nT,  Z  24.5 nT

Further spikes occurred between 00:30 and 02:15 UT on the following day 2025-04-28 with another smaller spike at  09:25 UT

Image
Average 10 min noise level  : 2025-04-27 20:00 to 23:00 (stable, all 3 axes)  = 17.0 nT (1 SD),  Y 10.5 nT,  Z  22.7 nT

The above lines in above graph use 5 min moving averages and therefore to see the detail the spikes were graphed using raw 10s readings and are shown below.

Image

Image

Image

Y Axis Test (2025-04-29)

Begun a test at 09:19 (08:19 UT) using just Y & Z magnetometers to see if the spikes in Y readings still occur when the X axis magnetometer isn't powered

Average 10 min noise level  : 2025-04-29 09:00 to 12:00 (stable, 2 axes)  = X --  (1 SD),  Y 8.7 nT,  Z  17.5 nT

Image

 

Image

Image

 

 

Y Axis Remedial Plan (2025-04-18)

It seems that the Y Axis Magnetometer is compromised in some way (but needs to be confirmed with new tests on 2025-04-26).

Rather than dig up the existing 3-axis Sensor Fixture and pull out the working X & Z sensors and repair or install a new Y Axis magnetometer  it is proposed to build a dedicated new Y axis fixture for burial beside the existing fixture.

Requires
 - new FG-3+ magnetometer (£67.70 from https://robosavvy.co.uk/fgsensors-fg-3-sensor.html)
 - new fluxgate capacitors
      10 uF, 25v Tatalum Capacitor  (£ 1 + delivery)
     100 nF, 25v Ceramic Capacitor (£ 1 +delivery)
- new piping  (3 ends caps, 1 x 90 deg bend, 1 x 135 deg bend, 2 connectors - £10)
- new 15m length of Cat6 cable required (already available, but may opt for a shielded Cat6 cable given lessons / experience.

Work
  Assembly / test new magnetometer & fixture
  Part dig existing hole, install new Y Axis fixture
  Dig up trench,  thread new cablethrough corrugated pipe (if possible) or run cable beside corrugated piper

 | Review | SAM III | Site Prep | Assembly Design | Parts Assembly |  Testing | Final Assembly | Installation | Commissioning | Remedial Work |

Back to Top


Daytime Geomagnetic Storm (2025-04-16)

A major daytime geomagnetic storm occured today (2025-04-16) which was picked up by the Observatory's new Geomagnetic Station which is still undergoing commissioning tests & observations.

After performing a diagnostic test involving turning off power to the station's downhole temperature sensor at 13:23 UT (which produced a downward shift in readings from all 3 magnetometers (X, Y & Z) it was noticed at around 14:20 that the readings from Station's X and Z sensors had been progressively increasing and the reading from Y sensor began to fluctuate more than it had been.   

Graph of Magnetometer Observations (2025-04-16)

Image

A check-in to the latest Scandinavian Stacked Magnetogram Charts showed that there was significant ongoing geomagnetic activity, and there was a good correspondance between deflections is Scandinavian Magnetogram lines (H Component) and those seen in the Station's X Component plot.

Plot of Scandinavian Magnetometer Readings with an overlay of the X Component from Clair Observatory's geomagnetic station
24h chart beginning at 2025-04-15 18:00 UT

Image

Note: The shift in X readings at 13:23 UT in original data has been corrected in order to make the Clair Observatory's magnetogram.

NASA Auroral Forecast for 18:00 UT is shown below.

Image

Back to Top